OpenStudy (anonymous):

fx=x+10 find (fof-) (10)

OpenStudy (anonymous):

f(f-1(10))= fof-1(10) f-1(x) = x-10 f-1(10)= 0 f(0) = 10 fof-1(10) =10

OpenStudy (anonymous):

fx=x+10 find (fof-) (10), the prob may be f of -10 f(-10)=-10+10 =0